OpenStudy (kmullis6):

PLEASE PLEASE PLEASE HELP! I'M REALLY STRUGGLING WITH THIS AND I KNOW ITS NOT MATH, BUT NO ONE'S ANSWERING AND I CANT FIND IT IN MY LESSONS /: What properties of sound determine the volume of sound? Is this affected by the motion of the sound source?

OpenStudy (anonymous):

Volume of sound? There's no such thing.

OpenStudy (kmullis6):

@KSameer73 I copied it off my assignment /:

OpenStudy (anonymous):

Maybe it means "intensity".

OpenStudy (kmullis6):

The first question was ''What properties of sound determine the pitch of sound? How is this affected by motion of the sound source? Be specific.'' but I answered it already, so idk if that helps you understand the question a little more...

OpenStudy (mathmale):

I'd substitute the word "intensity" (or "loudness" for "volume.

OpenStudy (anonymous):

it is not affected by the motion of the sound source and your answer might be Aux if that's an option.

OpenStudy (mathmale):

Well, compare a woofer and a tweeter loud speaker. Which one do you think consumes the more electrical energy?

OpenStudy (kmullis6):

@Zariel<3sthelegomovie there's no options. /: it's a write out of 2-3 sentences type answer thing i have to do

OpenStudy (kmullis6):

@mathmale idk what a woofer and tweeter are sorry. like woofer is loud, and tweeter is lower?

OpenStudy (mathmale):

Actually, if a train is coming towards you and blows its horn, the sound of the horn will change as the distance between u and the train decreases; this is called the Doppler effect. But that pertains more to the frequency of the sound than it does to the volume. Still, as the train approaches you, it will sound increasingly louder. But if you're on the train, the sound won't change.

OpenStudy (anonymous):

Loudness is the rate of power delivery.\[loudness=\frac{dP}{dt}\]

OpenStudy (anonymous):

ok in that case.. Pitch is a common term we give for frequency. So changing the pitch means changing the frequency of sound. Sound, like all things, travels a distance L = velocity X time = vt. If we set L = wavelength, then the time to travel one wavelength is called the periiod T. And frequency f = 1/T, the reciprocal of the period. So when L is wavelength, we have L = vT = v/f. This means, for a fixed speed of sound v, the wave length of that sound decreases as its frequency, its pitch goes higher. Or stated in reverse, f = v/L, the frequency increases as the wave length decreases. Now what happens when the source of that sound is moving? If its moving toward us, then V > v is the apparent velocity of sound. That is, the apparent velocity is just v + U = V where U is the velocity of the source. And, from F = V/L > v/L = f, we see that F > f. In other words, when the source is coming at us, we hear a higher pitch F > f than when the source is standing still. And the reverse, with the source moving away from us, gives us a lower pitch because the relative velocity of sound is now less than the standing still velocity. That is v - U is the apparent velocity so the frequency f decreases. if this helps.

OpenStudy (mathmale):

the comical thing here, @kmullis6, is that I'm deaf and don't know! But I suspect that a tweeter is of higher frequency and lower POWER.

OpenStudy (mathmale):

wow! quite a good discussion!

OpenStudy (kmullis6):

Would this answer the question, because I was thinking about putting it... ''Yes, it is affected. When the object is coming closer to you, the volume changes.'' or something like that? Sorry if I'm not grasping what y'all are telling me. Science is my worst subject. ):

OpenStudy (anonymous):

and @kmullis6 here's a free medal!

OpenStudy (kmullis6):

@Zariel<3sthelegomovie thanks!

OpenStudy (kmullis6):

@KSameer73 @mathmale @Zariel<3sthelegomovie Would this answer the question, because I was thinking about putting it... ''Yes, it is affected. When the object is coming closer to you, the volume changes.'' or something like that? Sorry if I'm not grasping what y'all are telling me. Science is my worst subject. ):

OpenStudy (anonymous):

yes when the object is coming closer to you, the volume changes thats the basics.

OpenStudy (anonymous):

Zariel is right.